In a shipment of 10 computers, 3 are defective. Three computers are randomly selected and tested. What is the probability that all three are defective if the first and second ones are not replaced after being tested?


1
Expert's answer
2022-05-27T13:52:10-0400

First one: probability=3/20

Second one: p= 2/19

third one: p= 1/18

So, the probability that all three are defective if the first and second ones are not replaced after being tested is 3/20*2/19*1/18 = 1/(3*20*19) =1/1140